OpenStudy (anonymous):

Two angles are supplementary and one of them is 4% less than 3 times the other

OpenStudy (anonymous):

angle 1 = a, angle 2 = b multiplying by 0.96 is the same as 4% less a = 0.96(3b) they are supplimentary meaning they add to 180 a + b = 180 solve for b b = 180 - a replace into other equation a = 0.96(3(180 - a) a = 0.96(540 - 3a) solve for a from here. replace answer into a + b = 180 to get angle b.

OpenStudy (anonymous):

two angles, assume one is x, the other one is 0.96*3*x=2.88x. so since these are supplementary x+2.88x=180, 3.88x=180, x=46.39, 133.61